OpenStudy (gabylovesyou):

A study of one thousand teens found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 10 hours. The population standard deviation is 2 hours. What is the 95% confidence interval for the mean? 9.88−10 hours 10−10.12 hours 9.88−10.12 hours 9.76−10.24 hours

OpenStudy (gabylovesyou):

@KendrickLamar2014

OpenStudy (debbieg):

What did you get for your answer?

OpenStudy (gabylovesyou):

C.. .-.

OpenStudy (debbieg):

Well done.

OpenStudy (gabylovesyou):

yay thanks.

OpenStudy (debbieg):

Sure thing. :) Just plug & chug in the formula. A confidence interval is just a "range" around a sample mean. Just use the formula to come up with what you add and subtract to the sample mean, and you are good to go. :)

OpenStudy (agent0smith):

\[\Large x \pm z* \frac{ \sigma }{ \sqrt n }\]z* for 95% is 1.96 \[\Large 10 \pm 1.96 \frac{ 2}{ \sqrt {1000} }\] So your answer looks right.